OpenStudy (anonymous):

A passenger in an airplane flying an altitude of 10 kilometers sees two towns directly to the left of the plane. The angles of depression to the towns are 28 degrees and 55 degrees. How far apart are the towns?

OpenStudy (anonymous):

I think if you plug in 10cos(55) into your calculator you will get one and the other should be 10cos(28)

OpenStudy (dumbcow):

 angle APB = 55-28 = 27 27+35 = 62 For triangle PAO, tan 62 = AO/10 --> AO = 10*tan(62) For triangle PBO, tan 35 = BO/10 --> BO = 10*tan(35) AB = AO -BO = 10tan(62) -10tan(35)